# Spontaneous Symmetry Breaking in Graviweak Theory

### v1

@@ III Graviweak Lagrangian in Two-Spinor Notation
We begin by reviewing the two-spinor (or $\mathbb{C}$)) notation. In four-dimensional spacetime, the Lorentz algebra $\mathfrak{so}(3,1)$ is locally isomorphic to $\mathfrak{sl}(2,\mathbb{C})\oplus\mathfrak{sl}(2,\mathbb{C})$, corresponding to the self-dual and anti-self-dual parts of the spin connection. Using the Infeld–van der Waerden map Jacobson:1988fb ; Penrose:1984mf ; Capovilla:1991kp , spacetime indices $a=0,1,2,3$ are replaced by spinor indices via the soldering form $\sigma^{a}_{AA^{\prime}}$:
@@
An antisymmetric tensor such as the spin connection $\omega^{ab}$ decomposes into self-dual ()(-)( - ) and anti-self-dual (+)(+)( + ) parts as per (2), with:
@@
where $A^{AB}=A^{(AB)}$ is the right-chiral (SD) $\mathbb{C}$) connection, and $\bar{A}^{A^{\prime}B^{\prime}}=\bar{A}^{(A^{\prime}B^{\prime})}$ is its left-chiral (ASD) counterpart. The tetrad $e^{a}$ is similarly rewritten as:
@@
$\mathcal{L}_{\text{YM}}=-\frac{i}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}},$ (22)
@@
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}},$ (23)
where $D^{(+)}$ is the spinor covariant derivative acting on left-handed spinor $\psi_{A^{\prime}}$ via the connection $\bar{A}^{A^{\prime}}_{\,\,\,\,B^{\prime}}$ and
@@
$\mathcal{L}=-\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{B}_{\,\,\,A^{\prime}}\wedge F_{AB}-\frac{i}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}}+\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}.$ (25)
@@ IV.1 Structure of the Gauge Fields
$\mathfrak{so}(1,3+N)\cong\mathfrak{so}(1,3)\oplus\mathfrak{so}(N).$ (26)
The generators $M_{AB}$ obey the commutation relation $[M_{AB},M_{CD}]=\eta_{AD}M_{BC}-\eta_{AC}M_{BD}-\eta_{BD}M_{AC}+\eta_{BC}M_{AD}$ and have indices running as $A,B,C...=a,b,c,...,i,j,k,...=0,...,3+N$, where $a,b,c,..=0,1,2,3$ are in $\mathfrak{so}(1,3)$ and $i,j,k,...=4,...,3+N$ in $\mathfrak{so}(N)$. We therefore have $\frac{(N+4)(N+3)}{2}$ generators, which break into the 6 generators of $SO(1,3)$ plus the $\frac{N(N-1)}{2}$ generators of $SO(N)$.
The gauge field is valued in the Lie algebra of $\mathfrak{so}(1,3+N)$, with gauge field components $\mathcal{A}^{A}$ and its field strength is given by:
@@ IV.2 The $\mathbb{SO(1,6)}$ case
$S$ $=\int\epsilon_{A_{1}...A_{7}}\mathcal{F}^{A_{1}A_{2}}\wedge\mathcal{F}^{A_{3}A_{4}}V^{A_{5}A_{6}A_{7}}$ (31)
$+\mathcal{F}^{AB}\wedge*\mathcal{F}_{AB}$
where we have introduced the auxiliary 0-form fields $V^{ABC}$. The second term corresponds to the kinetic term of the gauge field with the spacetime hodge dual; the first term can be written explicitly in terms of its $\mathfrak{so}(1,3)$ and $\mathfrak{so}(3)$ components as
$S_{\text{int}}$ $=S_{4,0}+S_{3,1}+S_{2,2}^{(a)}+S_{2,2}^{(b)}+S_{1,3},$ (32)
with
$S_{4,0}$ $=\int\;\epsilon_{abcd\,ijk}\;(\mathcal{F}^{ab}\wedge\mathcal{F}^{cd})\;V^{ijk},$ (34)
$S_{3,1}$ $=-12\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ab}\wedge\mathcal{F}^{ci}\bigr{)}\;V^{djk},$ (35)
$S_{2,2}^{(a)}$ $=6\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ab}\wedge\mathcal{F}^{ij}\bigr{)}\;V^{cd\,k},$ (36)
$S_{2,2}^{(b)}$ $=-12\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ai}\wedge\mathcal{F}^{bj}\bigr{)}\;V^{cd\,k},$ (37)
$S_{1,3}$ $=-4\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ai}\wedge\mathcal{F}^{jk}\bigr{)}\;V^{bcd}.$ (38)
At this stage, one can consider breaking the symmetry in order to recover the Einstein-Cartan gravity sector along with the new interaction Einstein-Cartan/Yang-Mills interaction term. The first one arises from $S_{4,0}$ with the following symmetry breaking ansatz:
@@
$\mathcal{F}^{a}_{~{}i}=\frac{1}{\ell}\Bigl{(}T^{a}\,n_{i}+e^{a}\wedge n_{j}A^{j}{}_{i}\Bigr{)},$ (42)
where the first equation has the same form as the one in MacDowell-Mansouri theories MacDowell:1977jt . We need an additional condition on $V^{ABC}$ in order to eliminate the remaining terms in the action. This can be achieved with:
$V^{ABC}=\varphi\epsilon^{ijk}\delta^{A}_{i}\delta^{B}_{j}\delta^{c}_{k}+\delta^{A}_{a}\delta^{B}_{b}\delta^{C}_{k}V^{abk}.$ (43)
Implementing the above ansatz for $\phi$ and $V$, action $S_{4,0}$ yields the Einstein-Cartan action plus the Euler term:
$S_{EC}$ $=6\int\varphi~{}\epsilon_{abcd}(R^{ab}+\ell^{-2}e^{a}\wedge e^{b})$ (44)
$\wedge(R^{cd}+\ell^{-2}e^{c}\wedge e^{d}),$
with the identifications $\Lambda=-3/\ell^{2}$ and $\varphi=\frac{1}{16\pi G}\frac{\ell^{2}}{24}$.
The second result of this broken symmetry is the Einstein-Cartan-Yang-Mills coupling that stems from action $S_{2,2}^{(a)}$ which takes the form:
$S_{RF}$ $=6\int\;\epsilon_{abcd}(R^{ab}+\ell^{-2}e^{a}\wedge e^{b})\wedge\epsilon_{ijk}F^{ij}V^{cd\,k}.$ (45)
The full action after symmetry breaking can be recasted as:
$S$ $=\frac{1}{16\pi G}\int e^{a}\wedge e^{b}\wedge\star R_{ab}-\frac{3}{2\Lambda}R^{ab}\wedge\star R_{ab}$ (46)
$-\frac{\Lambda}{12}\epsilon_{abcd}e^{a}\wedge e^{b}\wedge e^{c}\wedge e^{d}$
$+6\int\;\epsilon_{abcd}(R^{ab}-\frac{\Lambda}{3}e^{a}\wedge e^{b})\wedge\epsilon_{ijk}F^{ij}V^{cd\,k}$
$+\mathcal{F}^{AB}\wedge*\mathcal{F}_{AB}.$
@@ V Conclusions and Outlook
Further work could explore detailed phenomenological models to quantify precisely how graviweak-induced parity violations could manifest observationally. Additionally, theoretical investigations could seek a deeper understanding of the interplay between graviweak symmetry breaking and established cosmological models, particularly in inflationary and pre-inflationary epochs. Exploring connections with quantum gravity approaches such as loop quantum gravity Rovelli:2004tv ; Alexander:2011qfa might also provide valuable insights into the fundamental structure of space-time and particle interactions. Finally, computational simulations of the early universe that incorporate graviweak effects could greatly enhance our predictive power and help guide future experimental efforts.
@@ Acknowledgements.
We thank Aaron Hui and Jacob Kuntzleman for helpful conversations. BA was supported by FCT Grant No. 2021.05694.BD and JM was partly supported by STFC Consolidated Grant ST/T000791/1.
@@ Appendix A Translation into the two-spinor notation
$\frac{1}{2}\epsilon_{abcd}\Sigma^{ab}\wedge R^{cd}=i(\Sigma^{(-)ab}\wedge R^{(-)}_{ab}-\Sigma^{(+)ab}\wedge R^{(+)}_{ab}).$ (47)
@@
$\mathcal{L}_{\text{YM}}=\frac{1}{4g_{+}^{2}}R^{(+)\,ab}\wedge\star R^{(+)}_{ab}=-\frac{i}{2g_{+}^{2}}\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}}.$ (58)
@@
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\star e^{a}\wedge\bar{\psi}_{L}\gamma_{a}D^{(+)}\psi_{L}=\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{A}\sigma^{a\,AA^{\prime}}D^{(+)}\psi_{A^{\prime}}$ (59)
$=\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}$ (60)

### v2

@@ III Graviweak Lagrangian in Two-Spinor Notation
[…]
@@
An antisymmetric tensor such as the spin connection $\omega^{ab}$ decomposes into self-dual ()(-) and anti-self-dual $(+)$ parts as per (2), with:
@@
[…]
@@
$\mathcal{L}_{\text{YM}}=-\frac{1}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\widetilde{\bar{F}}_{A^{\prime}B^{\prime}},$ (23)
@@
$\mathcal{L}_{\text{fermion}}=-\frac{i}{2}\,\star e^{AA^{\prime}}\wedge[\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}+\bar{\psi}_{A^{\prime}}\,D^{(-)}\psi_{A}],$ (24)
where $D^{(\pm)}$ is the spinor covariant derivative acting on left- and right-handed spinors $\psi_{A^{\prime}}$ and $\psi_{A}$ via the connections $\bar{A}^{A^{\prime}}_{\,\,\,\,B^{\prime}}$ and $A^{A}_{\,\,\,\,B}$. We also have
@@
$\mathcal{L}=-\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{B}_{\,\,\,A^{\prime}}\wedge F_{AB}-\frac{1}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\widetilde{\bar{F}}_{A^{\prime}B^{\prime}}-\frac{i}{2}\,\star e^{AA^{\prime}}\wedge[\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}+\bar{\psi}_{A^{\prime}}\,D^{(-)}\psi_{A}].$ (26)
@@ IV.1 Structure of the Gauge Fields
$\mathfrak{so}(1,3+N)\;\simeq\;\mathfrak{so}(1,3)\oplus\mathfrak{so}(N)\oplus(\text{coset})$ (27)
The generators $M_{AB}$ obey the commutation relation $M_{AB},M_{CD}]=\eta_{AD}M_{BC}-\eta_{AC}M_{BD}-\eta_{BD}M_{AC}+\eta_{BC}M_{AD}$ and have indices running as A,B,C=a,b,c,,i,j,k,=0,,3+NA,B,C...=a,b,c,...,i,j,k,...=0,...,3+N, where $a,b,c,..=0,1,2,3$ are in $\mathfrak{so}(1,3)$ and i,j,k,=4,,3+Ni,j,k,...=4,...,3+N in $\mathfrak{so}(N)$. We therefore have $\frac{(N+4)(N+3)}{2}$ generators, which break into the 6 generators of $SO(1,3)$ plus the $\frac{N(N-1)}{2}$ generators of $SO(N)$, and the remaining ones are associated with the coset.
The gauge field is valued in the Lie algebra of $\mathfrak{so}(1,3+N)$, with gauge field components $\mathcal{A}^{AB}$ and its field strength is given by:
@@ IV.2 The $\mathbb{SO(1,6)}$ case
$S$ $=\int\frac{\alpha}{2}\epsilon_{A_{1}...A_{7}}\mathcal{F}^{A_{1}A_{2}}\wedge\mathcal{F}^{A_{3}A_{4}}V^{A_{5}A_{6}A_{7}}$ (32)
$+\int\alpha\mathcal{F}^{AB}\wedge\widetilde{\mathcal{F}}_{AB}+\frac{\lambda}{4}(V_{ABC}V^{ABC}-1)^{2}\widetilde{1}$,
where we have introduced an $SO(1,6)$ Higgs field $V^{ABC}$ with the associated Mexican hat potential, and $\alpha$ and $\lambda$ are coupling constants. In the unitary gauge we take a vev $V_{0}^{ijk}=\frac{1}{3!}\epsilon^{ijk}$ (so $V_{0}\cdot V_{0}=1$), which spontaneously breaks $SO(1,6)$ into the internal $SO(3)$ block and the Lorentz $SO(1,3)$ block. Here $\epsilon^{ijk}$ is the Levi–Civita tensor of the internal $SO(3)$.
At this stage, one can consider breaking the symmetry in order to recover the Einstein-Cartan gravity sector along with the new interaction Einstein-Cartan/Yang-Mills interaction term. We follow a MacDowell-Mansouri-like symmetry breaking ansatz:
@@
$\mathcal{F}^{a}_{~i}=\frac{1}{\ell}\Bigl(T^{a}\,n_{i}+e^{a}\wedge P_{i}\Bigr),$ (36)
where $P_{i}=dn^{i}+n_{j}A^{j}{}_{i}$, $T^{a}\equiv De^{a}=de^{a}+\omega^{a}{}_{b}\wedge e^{b}$ is the torsion, and the first equation has the same form as the one in MacDowell-Mansouri theories MacDowell:1977jt . When the field $V$ condensates at its vev, the first term in (32) becomes
$\epsilon_{A_{1}...A_{7}}\mathcal{F}^{A_{1}A_{2}}\wedge{\mathcal{F}}^{A_{3}A_{4}}V^{A_{5}A_{6}A_{7}}=\epsilon_{abcd}\mathcal{F}^{ab}\wedge\mathcal{F}^{cd},$
more specifically:
$\frac{1}{2}\epsilon_{abcd}\mathcal{F}^{ab}\wedge\mathcal{F}^{cd}=\frac{1}{2}\epsilon_{abcd}R^{ab}\wedge R^{cd}+\frac{1}{\ell^{4}}\Sigma^{ab}\wedge\widetilde{\Sigma}_{ab}-\frac{2}{\ell^{2}}R^{ab}\wedge\widetilde{\Sigma}_{ab},$ (37)
where we defined the area 22-form $\Sigma^{ab}=e^{a}\wedge e^{b}$. The second term in (32) corresponds to the kinetic term of the gauge field with the spacetime hodge dual and it can be written explicitly in terms of its $\mathfrak{so}(1,3)$ and $\mathfrak{so}(3)$ components as
$\mathcal{F}^{AB}\wedge\widetilde{\mathcal{F}}_{AB}=\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}_{ab}+2\mathcal{F}^{ai}\wedge\widetilde{\mathcal{F}}_{ai}+$
$+\mathcal{F}^{ij}\wedge\widetilde{\mathcal{F}}_{ij}.$ (38)
Using equations (34) to (36), the terms above can be explicitly written as
$\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}_{ab}=R^{ab}\wedge\widetilde{R}_{ab}+\frac{1}{\ell^{4}}\Sigma^{ab}\wedge\widetilde{\Sigma}_{ab}-\frac{2}{\ell^{2}}R^{ab}\wedge\widetilde{\Sigma}_{ab},$
and
$\mathcal{F}^{ai}\wedge\widetilde{\mathcal{F}}_{ai}=\frac{1}{\ell^{2}}T^{a}\wedge\widetilde{T}_{a}+\frac{3}{\ell^{2}}P^{i}\wedge\widetilde{P}_{i},$ (39)
where we use the fact that $n_{i}P^{i}=0$.
The expansion around the vev, $V=V_{0}+\delta V$, leads to very interesting interactions between gravity and the weak force, but their detailed analysis is beyond the scope of this work. These terms are of the form
$\epsilon_{abcd\,ijk}\;(\mathcal{F}^{ab}\wedge\mathcal{F}^{cd})\;\delta V^{ijk},$ (40)
$\epsilon_{abcd\,ijk}\;\bigl(\mathcal{F}^{ab}\wedge\mathcal{F}^{ci}\bigr)\;\delta V^{djk},$ (41)
$\epsilon_{abcd\,ijk}\;\bigl(\mathcal{F}^{ab}\wedge\mathcal{F}^{ij}\bigr)\;\delta V^{cd\,k},$ (42)
$\epsilon_{abcd\,ijk}\;\bigl(\mathcal{F}^{ai}\wedge\mathcal{F}^{bj}\bigr)\;\delta V^{cd\,k},$ (43)
$\epsilon_{abcd\,ijk}\;\bigl(\mathcal{F}^{ai}\wedge\mathcal{F}^{jk}\bigr)\;\delta V^{bcd}.$ (44)
Among these, the term of utmost interest is the direct coupling between the gravitational curvature and the YM field strength of the weak force:
$\mathcal{L}_{\text{RF}}=\alpha\epsilon_{abcd}(R^{ab}-\ell^{-2}e^{a}\wedge e^{b})\wedge\epsilon_{ijk}F^{ij}\delta V^{cd\,k}.$ (45)
The first two terms in the action (32) can then combine to give
$\frac{1}{2}\epsilon_{abcd}\mathcal{F}^{ab}\wedge\mathcal{F}^{cd}+\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}_{ab}$ $=$ $\frac{1}{2}\epsilon_{abcd}R^{ab}\wedge R^{cd}+R^{ab}\wedge\widetilde{R}_{ab}+\frac{2}{\ell^{4}}\Sigma^{ab}\wedge\widetilde{\Sigma}_{ab}-\frac{4}{\ell^{2}}R^{ab}\wedge\widetilde{\Sigma}_{ab}$ (46)
$+F^{ij}\wedge\widetilde{F}_{ij}+\frac{1}{\ell^{2}}T^{a}\wedge\widetilde{T}_{a}+\frac{3}{\ell^{2}}P^{i}\wedge\widetilde{P}_{i}.$
Choosing the constants $\alpha$ and $\ell$ so that we can recover EC with cosmological constant $\Lambda$:
$\alpha$ $=$ $-\frac{1}{64\pi G}\,\frac{3}{\Lambda}\,,\qquad\frac{1}{\ell^{2}}\;=\;\frac{\Lambda}{3}\,$, (47)
leads to the final Lagrangian written as:
$\mathcal{L}$ $=$ $\frac{1}{32\pi G}\epsilon_{abcd}\left(R^{ab}-\frac{\Lambda}{6}e^{a}\wedge e^{b}\right)\wedge e^{c}\wedge e^{d}-\frac{1}{64\pi G}\left(T^{a}\wedge\widetilde{T}_{a}+3P^{i}\wedge\widetilde{P}_{i}\right)$ (48)
$-\frac{3}{64\pi G\Lambda}\left(F^{ij}\wedge\widetilde{F}_{ij}+R^{ab}\wedge\widetilde{R}_{ab}+\frac{1}{2}\epsilon_{abcd}R^{ab}\wedge R^{cd}\right)+\mathcal{L}_{\text{RF}},$
@@ V Conclusions and Outlook
Much work remains to be done. Further studies could explore detailed phenomenological models to quantify precisely how graviweak-induced parity violations could manifest observationally. The remarkable coincidence between the frequency sensitivity window of the upcoming LISA mission LISA:2024hlh and the electroweak energy scale (see for example Caprini:2018mtu ) has not escaped us. A non-minimal coupling between gravitons and the weak gauge bosons would have an impact on the stochastic gravitational wave background. Could our model induce a resonance in the production of gravity waves at the electroweak scale? And would its parity violating features make the corresponding spike inimitable via other mechanisms? Additionally, theoretical investigations could seek a deeper understanding of the interplay between graviweak symmetry breaking and established cosmological models, particularly in inflationary and pre-inflationary epochs. Exploring connections with quantum gravity approaches such as loop quantum gravity Rovelli:2004tv ; Alexander:2011qfa might also provide valuable insights into the fundamental structure of space-time and particle interactions. A variation of this model followup , for example, would connect directly with the work on area metrics emerging from spin foams Borissova:2022clg ; Borissova:2023yxs .
@@ Acknowledgements.
We thank Chiara Caprini, Aaron Hui, and Jacob Kuntzleman for helpful conversations. BA was supported by FCT Grant No. 2021.05694.BD and JM was partly supported by STFC Consolidated Grant ST/T000791/1.
@@ Appendix A Translation into the two-spinor notation
[…]
@@
$\mathcal{L}_{\text{YM}}=-\frac{1}{4g^{2}}R^{(+)\,ab}\wedge\widetilde{R}^{(+)}_{ab}=-\frac{1}{2g^{2}}\bar{F}^{A^{\prime}B^{\prime}}\wedge\widetilde{\bar{F}}_{A^{\prime}B^{\prime}}.$ (60)
@@
$\mathcal{L}_{\text{fermion}}=-\frac{i}{2}\star e^{a}\wedge\bar{\psi}_{L}\gamma_{a}D^{(+)}\psi_{L}=-\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{A}\sigma^{a\,AA^{\prime}}D^{(+)}\psi_{A^{\prime}}$ (61)
$=-\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}$ (62)
@@
Appendix B
An alternative action that accomplishes the spontaneous symmetry breaking needed to remove the left-handed Einstein-Hilbert term and the right-handed quadratic term is the following:
$S$ $=$ $\frac{\alpha}{2}\int\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}_{ab}+\frac{i\alpha}{4}\int\frac{\phi}{v}\,\epsilon_{abcd}\,\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}^{cd}+$ (65)
$+i\alpha\int\frac{\phi}{v}\,\mathcal{F}^{ab}\wedge\mathcal{F}_{ab}+\int V(\phi)\,\widetilde{1}\,$.
Here $\alpha$ is a coupling constant, $\phi$ is the pseudoscalar Higgs field and $V(\phi)=\frac{\lambda}{4}(\phi^{2}-v^{2})^{2}$ is the potential given above in section II. This action is also parity invariant before symmetry breaking. The first term is a kinetic term for the curvature, the second is an $SO(1,3)$ invariant term combining both Hodge duals in spacetime and in group indices, and the third is a Pontryagin-like term. Upon symmetry breaking of the pseudoscalar, we can simplify this new action as follows. Further using (34) leads to
$\frac{1}{2}\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}_{ab}+\frac{i}{4}\epsilon_{abcd}\mathcal{F}^{ab}\wedge\widetilde{\mathcal{F}}^{cd}=\mathcal{F}^{(+)ab}\wedge\widetilde{\mathcal{F}}^{(+)}_{ab}$=
$=R^{(+)ab}\wedge\widetilde{R}^{(+)}_{ab}-\frac{2}{\ell^{2}}\,\Sigma^{(+)}_{ab}\wedge\widetilde{R}^{(+)ab}+\frac{1}{\ell^{4}}\Sigma^{(+)}_{ab}\,\wedge\widetilde{\Sigma}^{(+)ab},$ (66)
Putting these together, equation (66) takes the form
[…]
Let us expand each term in the above equation in terms of its constituents as per the MacDowell-Mansouri formalism.
$\mathcal{F}^{ab}\wedge\mathcal{F}_{ab}=R^{ab}\wedge R_{ab}-\frac{2}{\ell^{2}}\,e_{a}\wedge e_{b}\wedge R^{ab}=$
$=R^{ab}\wedge R_{ab}-\frac{2}{\ell^{2}}R^{(+)ab}\wedge\Sigma^{(+)}_{ab}-\frac{2}{\ell^{2}}R^{(-)ab}\wedge\Sigma^{(-)}_{ab}$.
After spontaneous symmetry breaking and the pseudoscalar field takes the value $\phi=v$, then the Lagrangian takes the form
$\mathcal{L}$ $=$ $\frac{-2i\alpha}{\ell^{2}}\Sigma_{ab}^{(-)}\wedge R^{(-)ab}+\frac{i\alpha}{\ell^{4}}\Sigma^{(-)ab}\wedge\Sigma^{(-)}_{ab}$+ (68)
$+{\alpha}R_{ab}^{(+)}\wedge\widetilde{R}^{(+)ab}+i\alpha R^{ab}\wedge R_{ab}.$
If we set
$\alpha$ $=$ $-\frac{1}{32\pi G}\,\frac{3}{2\Lambda}\,,\qquad\frac{1}{\ell^{2}}\;=\;\frac{2\Lambda}{3}\,$, (69)
then the Lagrangian takes the form similar to that shown in section II:
$\mathcal{L}$ $=$ $\frac{1}{32\pi G}\epsilon_{abcd}\left(e^{a}\wedge e^{b}\wedge R^{(-)cd}-\frac{\Lambda}{6}e^{a}\wedge e^{b}\wedge e^{c}\wedge e^{d}\right)+$ (70)
$-\frac{3}{64\pi G\Lambda}R_{ab}^{(+)}\wedge\widetilde{R}^{(+)ab}.$
Note that the gauge coupling in this model is controlled by the value of the gravitational constant and the cosmological constant.